OpenStudy (anonymous):

how to factor out by grouping 15x^2-34x+15

OpenStudy (anonymous):

Traditionally, grouping is used when the expression has a degree of 3.

OpenStudy (experimentx):

one way of doing: find two number such that the product is 15x15 and sum is -34

OpenStudy (anonymous):

BUT. :) I got it. When you group, you need 4 terms to work with. So you need to split the middle term (-34x) into two terms: 15x^2 - 9x -25x + 15 Then group: (15x^2 -9x) + (-25x+15) Factor: 3x(5x-3) -5(5x-3) Then finalize: (3x-5)(5x-3) :) phew almost didn't get that one

OpenStudy (anonymous):

could you tell me how you splitted the term of -34 wombat

OpenStudy (anonymous):

-34 = -9-25 Pretty much, I looked at the terms 15x^2 and 15 These numbers need something divisible by 3 or 5 (or in rare cases, 15) So i looked for two numbers that add/subtract to get -34 that contained factors of 3 and/or 5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

7x^2-19x-6=0

OpenStudy (anonymous):

is this a new question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

try 7x^2 - 14x - 3x -6 =0 and proceed the same way i did above :)

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

sure :)